An RLC circuit has a 65.0 Ω resistor, a 7.80 mH inductor and a 230. nF capacitor. What is the peak current if the voltage source oscillates with a peak voltage of 7.10 V, with a frequency of 4.79×103 Hz?


magnitude:
Answer #1

We know,

I_{o} = \frac{V_{o}}{Z}

where,

Z=\sqrt{R^2 + (X_L -X_C)^2}

So, Computing Z:

Z=\sqrt{65^2 + (2*\pi*f_o*L -\frac{1}{2*\pi*f_o*C})^2}

Z=\sqrt{4225 + (2*\pi*4.79*10^3*7.8*10^{-3} -\frac{1}{2*\pi*4.79*10^3*230*10^{-9}})^2}

Z=\sqrt{4225 + (2*\pi*4.79*7.8 -\frac{10^6}{2*\pi*4.79*230})^2}

Z=\sqrt{4225 + (234.63336 -1.445*10^{-4}*10^6)^2}

Z=\sqrt{4225 + (234.63336 -144.5)^2}

Z=\sqrt{4225 + 8124.02}

Z=111.126\ \Omega

Now,

I_o = \frac{7.1}{111.126} = 0.0639\ A = 63.9\ mA

PLEASE RATE IF YOU LIKE IT